# Orbital Depot Location Optimization for Satellite Constellation Servicing with Low-Thrust Transfers

### v1

@@ Abstract
$i$/$j$  = index representing a satellite
@@ 2.1 Continuous Location-Routing Problem for Orbital Depots
Figure 1 illustrates a typical example of a solution for the CLRPOD. The scenario involves six satellites serviced by servicer spacecrafts from two depots on the initial orbits (Fig. 1a). Figure 1b shows the optimal solution, including the servicing sequences (or routes) and the final orbits of the depots. At Depot 1, the servicer operates along a single route. In contrast, at Depot 2, the servicer initially visits two satellites, returns to Depot 2 to resupply, and then services the remaining two satellites. For illustrative purposes, transfers between orbits are depicted as simple arrows, while actual trajectories follow optimized low-thrust transfer paths. The orbital elements of Depots 1 and 2 (i.e., semimajor axis, inclination, right ascension of the ascending node, eccentricity, and argument of perigee) are also optimized for minimizing the propellant EMLEOs. For simplicity, this study assumes circular orbits, excluding eccentricity and perigee argument from the optimization process. Nonetheless, the methodology presented in this paper can be extended to general elliptical orbits.
@@ 2.2 Low-Thrust Transfer Between Circular and Tilted Circular Orbits
where $R_{z}(-\Omega_{1})$/$R_{z}(-\Omega_{2})$, as well as $R_{x}(-i_{1})$/$R_{x}(-i_{2})$ are the rotation matrices about the z-axis and x-axis with angle $-\Omega_{1}$/$-\Omega_{2}$ and $-i_{1}$/$-i_{2}$, respectively. Note that $\Omega_{1}$ and $\Omega_{2}$ represent the right ascension ascending node (RAAN), and $i_{1}$ and $i_{2}$ denote the inclination of orbit 1 and orbit 2, respectively.
@@
$\Delta V=\sqrt{V_{1}^{2}-2V_{1}V_{2}\cos\left[\frac{\pi}{2}\Delta\theta\right]+V_{2}^{2}}$ (6)
where $V_{1}$ and $V_{2}$ are the circular orbit velocities for orbits 1 and 2, respectively, defined as $\sqrt{\mu/r_{1}}$ and $\sqrt{\mu/r_{2}}$, where $r_{1}$ and $r_{2}$ are the radii of circular orbits 1 and 2. Note that Eq. (6) is valid only for $0\leq\Delta\theta\leq 2$. If $\Delta\theta>2$, then $\Delta V=V_{1}+V_{2}$ [25]. So, Eq. (6) can be rewritten as follows.
@@ 2.3 Depot Launch and Deployment Costs
Accurate estimation of launch and insertion costs for orbital depots requires the consideration of both their masses and orbital elements, not merely summing depot masses. For example, if two depots have the same mass but are positioned at different altitudes, the one at a higher altitude incurs a greater launch cost. To address this challenge, we use the concept of EMLEO, which assumes each depot is transferred from a virtual circular orbit with radius $r_{0}$ to its intended orbit via a coplanar Hohmann transfer. The EMLEO is determined by adding the propellant mass required for the transfer. Additionally, we assume that the initial (hypothetical) burn is performed by the launch vehicle while the orbital depot executes the second burn. Mathematically, let $r_{0}$ be the EMLEO’s radius and $a_{k}$ be the orbital depot $k$’s radius. Then, the required $\Delta V$ by the launch vehicle ($\Delta V^{l}$) and depot’s deployment ($\Delta V^{d}$) are calculated as follows.
@@ 2.4 Mathematical Formulation of Orbital Depot Location-Routing Problem
$D_{e}$ =$=\{n_{d}n_{v},...,n_{d}(n_{v}+1)-1\}$ (14)
$T$ =$=\{n_{d}(n_{v}+1),...,n_{d}(n_{v}+1)+n_{t}-1\}$ (15)
where $n_{d}$ is the number of depots, $n_{v}$ is the number of available vehicles (routes) for each depot, and $n_{t}$ is the number of satellites. Note that the vehicles associated with each depot can be modeled as multiple depots, including the original depot and its duplications (virtual depots), for mathematical convenience. For example, if $n_{v}=1$, only original depots exist, meaning each depot has only a single route available. The set of original and virtual depots corresponding to an original depot $k\in D_{0}$ is defined as follows:
$D_{s}^{k}=\{k,k+n_{d},...,k+n_{d}(n_{v}-1)\}\qquad\forall k\in D_{0}$ (16)
@@
$\min_{{\bf\it z},{\bf\it u},{\bf\it\eta}}\left[\sum_{k\in D_{s}}\left(u_{k}-m_{k^{\prime}}^{\text{dry},s}\sum_{j\in T}z_{kj}^{k}\right)\phi_{k^{\prime}}({\bf\it\eta}_{k^{\prime}})\right]$ (17)
@@
$z_{kj}^{k}-z_{jk^{\prime\prime}}^{k}+\sum_{i\in T,i\neq j}\left(z_{ij}^{k}-z_{ji}^{k}\right)=0\qquad\forall k\in D_{s},j\in T$ (20)
$u_{k}+M\left(1-z_{kj}^{k}\right)\geq\left(u_{j}+m^{L}_{j}\right)\exp\left(\frac{\Delta V_{k^{\prime}j}({\bf\it\eta}_{k^{\prime}})}{g_{0}I_{sp}^{s}}\right)\qquad\forall k\in D_{s},j\in T$ (21)
@@
$u_{i}+M\left(1-z_{ik^{\prime\prime}}^{k}\right)\geq m^{\text{dry},s}_{k^{\prime}}\exp\left(\frac{\Delta V_{ik^{\prime}}({\bf\it\eta}_{k^{\prime}})}{g_{0}I_{sp}^{s}}\right)\qquad\forall k\in D_{s},i\in T$ (23)
$\left[\sum_{p\in D_{s}^{k}}\left(u_{p}-m_{k}^{\text{dry},s}\sum_{j\in T}z_{pj}^{p}\right)+m_{k}^{\text{dry},s}+m_{k}^{\text{dry},d}\right]\phi_{k}({\bf\it\eta}_{k})\leq m_{\text{max}}^{l}\qquad\forall k\in D_{0}$ (24)
@@
$z_{ik^{\prime\prime}}^{k}\in\{0,1\}\qquad\forall k\in D_{s},i\in T$ (27)
@@
Equation (17) is the objective function, which minimizes the overall propellant mass in terms of EMLEO. Index $k\in D_{s}$ represents either the original or virtual depot, while $k^{\prime}\in D_{0}$ is an original depot index for given $k\in D_{s}$ defined as $k^{\prime}(k)=\text{mod}(k,n_{d})$. $u_{k}$ is the mass before departure from depot $k$ (i.e., sum of the servicer’s dry mass, the required propellant mass, and the mass of all payloads to be serviced), $m_{k^{\prime}}^{\text{dry},s}$ is the dry mass of servicer $k^{\prime}$, $z_{ij}^{k}$ is the binary decision variable, which equals one if the vehicle $k$ travels from node $i$ to $j$ and zero otherwise, and $\phi_{k^{\prime}}$ is the EMLEO conversion factor to convert the propellant mass into EMLEO. ${\bf\it\eta}_{k^{\prime}}=[a_{k^{\prime}},i_{k^{\prime}},\Omega_{k^{\prime}}]$ is the set of orbital elements of depot $k^{\prime}$. Since $\phi_{k^{\prime}}$ is a nonlinear function of the semimajor axis of depot $k^{\prime}$ ($a_{k^{\prime}}$), the objective function is nonlinear. Equation (18) indicates that every start node (original or virtual depot) should be used at most once. Equation (19) enforces every given satellite should be visited exactly once, and Eq. (20) represents the flow continuity constraint. Equations (2123) ensure that the mass at each node reflects the transfer propellant mass calculated by the rocket equations. $u_{i}$ is the mass of servicer at node $i$ after deploy the payload ($m_{i}^{L}$), $M$ is a sufficiently large positive constant (big-M), $\Delta V_{ij}$ is the required velocity increment for orbit transfer from node $i$ to node $j$, and $I_{sp}^{s}$ is the specific impulse of the servicer. Equation (24) restricts the maximum launch mass for each depot in terms of the EMLEO. Note that $\Delta V_{k^{\prime}j}$, $\Delta V_{ik^{\prime}}$, and $\phi_{k}$ are the functions of ${\bf\it\eta}_{k^{\prime}}$ or ${\bf\it\eta}_{k}$, which make constraints nonlinear. Equations (2527) and Eq. (28) impose the binary and nonnegative constraints for the decision variables. Note that $k^{\prime\prime}\in D_{e}$ is defined as $k^{\prime\prime}(k)=n_{d}n_{v}+\text{mod}(k,n_{d})~{}\forall k\in D_{s}$, indicating the end node corresponding to the start node $k\in D_{s}$.
@@ 3.1 Mixed-Integer Linear Programming with Fixed Depots’ Locations
Assume all orbital elements of depots are fixed. Then, $\phi_{k}$, $\Delta V_{k^{\prime}j}({\bf\it\eta}_{k}^{\prime})$, and $\Delta V_{ik^{\prime}}({\bf\it\eta}_{k}^{\prime})$ are all constants. So, the problem becomes an extension of the VRP, which can be formulated as follows:
@@
$\left[\sum_{p\in D_{s}^{k}}\left(u_{p}-m_{k}^{\text{dry},s}\sum_{j\in T}z_{pj}^{p}\right)+m_{k}^{\text{dry},s}+m_{k}^{\text{dry},d}\right]\phi_{k}\leq m_{\text{max}}^{l}\qquad\forall k\in D_{0}$ (32)
Since all $\Delta V_{ij}$ and $\phi_{k}$ are constants, the lower bound of $M$ is determined as follows:
$M$ $\geq\quantity(\frac{m_{\text{max}}^{l}}{\min_{k\in D_{0}}\phi_{k}}+\max_{i\in T}m_{i}^{L})\max_{\begin{subarray}{c}i\in D_{s}\cup T\\ j\in T\cup D_{e}\end{subarray}}\exp\quantity(\frac{\Delta V_{ij}}{g_{0}I_{sp}^{s}})$ (33)
@@ 3.2 Nonlinear Programming with Fixed Routing Solution
$m^{p}({\bf\it\eta})$ $=\left(\left(m^{\text{dry},s}\exp\left[\frac{\Delta V_{n+1}}{g_{0}I_{sp}^{s}}\right]+m_{n}^{L}\right)\exp\left[\frac{\Delta V_{n}}{g_{0}I_{sp}^{s}}\right]+\cdots+m_{1}^{L}\right)\exp\left[\frac{\Delta V_{1}}{g_{0}I_{sp}^{s}}\right]-m^{\text{dry},s}-\sum_{i=1}^{n}m^{L}_{i}$ (34)
@@
where $A=m^{\text{dry,s}}\exp\left[\frac{\sum_{i=2}^{n}\Delta V_{i}}{g_{0}I_{sp}^{s}}\right]$, $B=\sum_{i=1}^{n}m_{i}^{L}\exp\left[\frac{\sum_{j=2}^{i}\Delta V_{j}}{g_{0}I_{sp}^{s}}\right]$, and $C=m^{dry,s}+\sum_{i=1}^{n}m_{i}^{L}$, which are all constants.
@@
$\min_{{\bf\it\eta}}J({\bf\it\eta})=\min_{{\bf\it\eta}_{k}}\sum_{k\in D_{0}}\sum_{r\in R_{k}}m_{r,k}^{p}({\bf\it\eta}_{k})\phi_{k}({\bf\it\eta}_{k})$ (37)
@@
$r_{\text{min}}$ $\leq a_{k}\qquad\forall k\in D_{0}$ (38)
@@
The partial derivative of $J$ with respect to ${\bf\it\eta}_{k}$ is given as:
$\frac{\partial J}{\partial{\bf\it\eta}_{k}}=\sum_{r\in R_{k}}\left(\frac{\partial m_{r,k}^{p}}{\partial{\bf\it\eta}_{k}}\phi_{k}+m_{r,k}^{p}\frac{\partial\phi_{k}}{\partial{\bf\it\eta}_{k}}\right)\qquad\forall k\in D_{0}$ (39)
For notational convenience, the subscript $k$ is omitted without loss of generality because the derivative that will be derived below can be applied to any $k$, and they are independent of each other. Mathematically,
@@
[…]
@@
Now, consider the derivative of $m^{p}_{r,k}$ with respect to $a_{k}$. Similarly, for notational convenience, the subscript $r$ and $k$ are omitted without loss of generality because the derivative that will be derived below can be applied to any route $r$ of depot $k$. From Eq. (7), $\Delta V_{1}$ and $\Delta V_{n+1}$ are the function of depot’s orbital elements, ${\bf\it\eta}$, given as follows.
@@
$\partialderivative{\Delta V_{1}}{a}$ $=\frac{1}{2\Delta V_{1}}\quantity(-\frac{1}{a^{2}}+\frac{1}{\sqrt{a^{3}a_{1}}}\cos\quantity[\frac{\pi}{2}\min\quantity{\Delta\theta_{1},2}])$ (46)
@@
$\partialderivative{\Delta V_{n+1}}{a}$ $=\frac{1}{2\Delta V_{n+1}}\quantity(-\frac{1}{a^{2}}+\frac{1}{\sqrt{a^{3}a_{n+1}}}\cos\quantity[\frac{\pi}{2}\min\quantity{\Delta\theta_{n+1},2}])$ (49)
@@
$\partialderivative{\Delta\theta_{1}}{i}$ =$=\begin{cases}-\frac{1}{\sqrt{1-\cos^{2}\Delta\theta_{1}}}\quantity(\cos i\sin i_{1}\cos\quantity(\Omega-\Omega_{1})-\sin i\cos i_{1}),&\Delta\theta_{1}\leq 2\\ 0,&\text{otherwise}\end{cases}$ (52)
$\partialderivative{\Delta\theta_{1}}{\Omega}$ =$=\begin{cases}\frac{1}{\sqrt{1-\cos^{2}\Delta\theta_{1}}}\sin i\sin i_{1}\sin\quantity(\Omega-\Omega_{1}),&\Delta\theta_{1}\leq 2\\ 0,&\text{otherwise}\end{cases}$ (53)
$\partialderivative{\Delta\theta_{n+1}}{i}$ =$=\begin{cases}-\frac{1}{\sqrt{1-\cos^{2}\Delta\theta_{n+1}}}\quantity(\cos i\sin i_{n+1}\cos\quantity(\Omega-\Omega_{n+1})-\sin i\cos i_{n+1}),&\Delta\theta_{n+1}\leq 2\\ 0,&\text{otherwise}\end{cases}$ (54)
$\partialderivative{\Delta\theta_{n+1}}{\Omega}$ =$=\begin{cases}\frac{1}{\sqrt{1-\cos^{2}\Delta\theta_{n+1}}}\sin i\sin i_{n+1}\sin\quantity(\Omega-\Omega_{n+1}),&\Delta\theta_{n+1}\leq 2\\ 0,&\text{otherwise}\end{cases}$ (55)
Note that $i_{1}$/$i_{n+1}$ and $\Omega_{1}$/$\Omega_{n+1}$ represent the inclination and RAAN of the first and last satellite orbits, respectively. Therefore, the first derivatives of the objective function can be calculated analytically, ensuring fast computation and accuracy for the L-BFGS-B algorithm.

### v2

@@ Abstract
ii/$j$  = index representing a satellite
@@ 2.1 Continuous Location-Routing Problem for Orbital Depots
It is important to note that this study focuses exclusively on regular, non-urgent maintenance and refueling operations, where service schedules can be fully planned in advance. In other words, mission time is not a critical constraint in the proposed problem. This assumption enables low-thrust propulsion systems to reduce overall fuel consumption, which aligns with minimizing EMLEO. In contrast, time-critical servicing missions such as emergency response or dynamic tasks require time constraints in the optimization and may involve high-thrust propulsion or other rapid deployment methods.
Figure 1 illustrates a typical example of a solution for the CLRPOD. The scenario involves six satellites serviced by servicer spacecrafts from two depots on the initial orbits (Fig. 1a). Figure 1b shows the optimal solution, including the servicing sequences (or routes) and the final orbits of the depots. At Depot 1, the servicer operates along a single route. In contrast, at Depot 2, the servicer initially visits two satellites, returns to Depot 2 to resupply, and then services the remaining two satellites. For illustrative purposes, transfers between orbits are depicted as simple arrows, while actual trajectories follow optimized low-thrust transfer paths. The orbital elements of Depots 1 and 2 (i.e., semimajor axis, inclination, right ascension of the ascending node, eccentricity, and argument of perigee) are also optimized for minimizing the propellant EMLEOs. For simplicity, this study assumes circular orbits, excluding eccentricity and argument of perigee from the optimization process. Nonetheless, the methodology presented in this paper can be extended to general elliptical orbits.
@@ 2.2 Low-Thrust Transfer Between Circular and Tilted Circular Orbits
where $R_{z}(-\Omega_{1})$$R_{z}(-\Omega_{2})$, as well as $R_{x}(-i_{1})$$R_{x}(-i_{2})$ are the rotation matrices about the z-axis and x-axis with angle $\Omega_{1}$$\Omega_{2}$ and $i_{1}$$i_{2}$, respectively. Note that $\Omega_{1}$ and $\Omega_{2}$ represent the right ascension ascending node (RAAN), and $i_{1}$ and $i_{2}$ denote the inclination of orbit 1 and orbit 2, respectively.
@@
where $V_{1}$ and $V_{2}$ are the circular orbit velocities for orbits 1 and 2, respectively, defined as $\sqrt{\mu/r_{1}}$ and $\sqrt{\mu/r_{2}}$, where $r_{1}$ and $r_{2}$ are the radii of circular orbits 1 and 2. The full derivation of Eq. (6) is detailed in previous literature [23, 25]. The unit of the inclination change, $\Delta\theta$, is radians. The factor of $\frac{\pi}{2}$ appears because it connects the desired tilt angle (relative inclination) change ($\Delta\theta$) with the yaw steering of the spacecraft, reflecting the average effectiveness of the out-of-plane thrust component when applied continuously throughout the orbital plane change. Note that if $\Delta\theta>2$, then $\Delta V=V_{1}+V_{2}$ [25].
@@ 2.3 Depot Launch and Deployment Costs
Accurate estimation of launch and insertion costs for orbital depots requires the consideration of both their masses and orbital elements, not merely summing depot masses. For example, if two depots have the same mass but are positioned at different altitudes, the one at a higher altitude incurs a greater launch cost. To address this challenge, we use the concept of EMLEO, which assumes each depot is transferred from a virtual circular orbit with radius $r_{0}$ to its intended orbit via a coplanar Hohmann transfer. The EMLEO is determined by adding the propellant mass required for the transfer. Additionally, we assume that the initial (hypothetical) burn is performed by the launch vehicle while the orbital depot executes the second burn. Mathematically, let $r_{0}$ be the radius used in calculating EMLEO and $a_{k}$ be the orbital depot $k$’s radius. Then, the required $\Delta V$ by the launch vehicle ($\Delta V^{l}$) and depot’s deployment ($\Delta V^{d}$) are calculated as follows.
@@
Note that this paper does not model the launch cost from Earth’s surface to the parking orbit. The primary focus is optimizing depot placement and the servicing sequence among known satellites. Incorporating detailed launch vehicle trajectory design and deployment options (e.g., dedicated, rideshare, piggyback missions, launch vehicle selection, or launch site choice) is beyond the current scope and could be explored in future work. Moreover, the EMLEO metric includes the transportation cost from parking orbit to the final servicing orbit, which allows for a meaningful optimization of in-space operations and propellant usage.
@@ 2.4 Mathematical Formulation of Orbital Depot Location-Routing Problem
$D_{e}$ $=\{n_{d}n_{v},...,n_{d}(n_{v}+1)-1\}$ (13)
$T$ $=\{n_{d}(n_{v}+1),...,n_{d}(n_{v}+1)+n_{t}-1\}$ (14)
where $n_{d}$ is the number of depots, $n_{v}$ is the number of available vehicles (routes) for each depot, and $n_{t}$ is the number of satellites. Note that the vehicles associated with each depot can be modeled as multiple depots, including the original depot and its duplications (virtual depots), for mathematical convenience. For example, if $n_{v}=1$, only original depots exist, meaning each depot has only a single route available.
To manage the formulation involving virtual depots, we define two auxiliary indices for each start node $k\in D_{s}$. First, $k^{\prime}(k)\in D_{0}$ denotes the original depot associated with the start node $k$, computed as $k^{\prime}(k)\equiv\text{mod}(k,n_{d})$. Second, $k^{\prime\prime}(k)\in D_{e}$ represents the corresponding end node, defined as $k^{\prime\prime}(k)\equiv n_{d}n_{v}+\text{mod}(k,n_{d})$. Both $k^{\prime}(k)$ and $k^{\prime\prime}(k)$ are thus functions of the start node index $k\in D_{s}$. The set of original and virtual depots corresponding to an original depot $k^{\prime}\in D_{0}$ is defined as follows:
$D_{s}^{k^{\prime}}=\{k^{\prime},k^{\prime}+n_{d},...,k^{\prime}+n_{d}(n_{v}-1)\}\qquad\forall k^{\prime}\in D_{0}$ (15)
@@
$\min_{{\bf\it z},{\bf\it u},{\bf\it\eta}}\left[\sum_{k\in D_{s}}\left(u_{k}-m_{k^{\prime}(k)}^{\text{dry},s}\sum_{j\in T}z_{kj}^{k}\right)\phi_{k^{\prime}(k)}({\bf\it\eta}_{k^{\prime}(k)})\right]$ (16)
@@
$z_{kj}^{k}-z_{jk^{\prime\prime}(k)}^{k}+\sum_{i\in T,i\neq j}\left(z_{ij}^{k}-z_{ji}^{k}\right)=0\qquad\forall k\in D_{s},j\in T$ (19)
$u_{k}+M\left(1-z_{kj}^{k}\right)\geq\left(u_{j}+m^{L}_{j}\right)\exp\left(\frac{\Delta V_{k^{\prime}(k)j}({\bf\it\eta}_{k^{\prime}(k)})}{g_{0}I_{sp}^{s}}\right)\qquad\forall k\in D_{s},j\in T$ (20)
@@
$u_{i}+M\left(1-z_{ik^{\prime\prime}(k)}^{k}\right)\geq m^{\text{dry},s}_{k^{\prime}(k)}\exp\left(\frac{\Delta V_{ik^{\prime}(k)}({\bf\it\eta}_{k^{\prime}(k)})}{g_{0}I_{sp}^{s}}\right)\qquad\forall k\in D_{s},i\in T$ (22)
$\left[\sum_{p\in D_{s}^{k^{\prime}}}\left(u_{p}-m_{k^{\prime}}^{\text{dry},s}\sum_{j\in T}z_{pj}^{p}\right)+m_{k^{\prime}}^{\text{dry},s}+m_{k^{\prime}}^{\text{dry},d}\right]\phi_{k^{\prime}}({\bf\it\eta}_{k^{\prime}})\leq m_{\text{max}}^{l}\qquad\forall k^{\prime}\in D_{0}$ (23)
@@
$z_{ik^{\prime\prime}(k)}^{k}\in\{0,1\}\qquad\forall k\in D_{s},i\in T$ (26)
@@
Equation (16) is the objective function, which minimizes the overall propellant mass in terms of EMLEO. $u_{k}$ is the mass before departure from depot $k$ (i.e., sum of the servicer’s dry mass, the required propellant mass, and the mass of all payloads to be serviced), $m_{k^{\prime}}^{\text{dry},s}$ is the dry mass of servicer $k^{\prime}$, $z_{ij}^{k}$ is the binary decision variable, which equals one if the vehicle $k$ travels from node $i$ to $j$ and zero otherwise, and $\phi_{k^{\prime}}$ is the EMLEO conversion factor to convert the propellant mass into EMLEO. ${\bf\it\eta}_{k^{\prime}}=[a_{k^{\prime}},i_{k^{\prime}},\Omega_{k^{\prime}}]$ is the set of orbital elements of depot $k^{\prime}$. Since $\phi_{k^{\prime}}$ is a nonlinear function of the semimajor axis of depot $k^{\prime}$ ($a_{k^{\prime}}$), the objective function is nonlinear.
The use of the EMLEO conversion factor ensures that the objective function inherently accounts for the delivery cost of both depot deployment and servicing operations. The physical meaning of the objective is the effective propellant mass required to conduct all servicing missions and deliver them in their respective orbits. Importantly, the model does not assume that 1 kg of propellant used by a servicer is equivalent to 1 kg used by a launch vehicle or depot. Instead, each mission stage (i.e., depot delivery and satellite servicing) is modeled using its respective propulsion parameters and $\Delta V$ requirements. Consequently, the EMLEO conversion factor appropriately combines launch, deployment, and servicing fuel consumption into a single objective value.
Equation (17) indicates that every start node (original or virtual depot) should be used at most once. Equation (18) enforces every given satellite should be visited exactly once, and Eq. (19) represents the flow continuity constraint. Equations (2022) ensure that the mass at each node reflects the transfer propellant mass calculated by the rocket equations. $u_{i}$ is the mass of servicer at node $i$ after deploy the payload ($m_{i}^{L}$), $M$ is a sufficiently large positive constant (big-M), $\Delta V_{ij}$ is the required velocity increment for orbit transfer from node $i$ to node $j$, and $I_{sp}^{s}$ is the specific impulse of the servicer. Equation (23) restricts the maximum launch mass for each depot in terms of the EMLEO. Note that $\Delta V_{k^{\prime}j}$, $\Delta V_{ik^{\prime}}$, and $\phi_{k^{\prime}}$ are the functions of ${\bf\it\eta}_{k^{\prime}}$, which make constraints nonlinear. Equations (2426) and Eq. (27) impose the binary and nonnegative constraints for the decision variables. For notational convenience, we will write $k^{\prime}(k)$ as $k^{\prime}$ and $k^{\prime\prime}(k)$ as $k^{\prime\prime}$ for the rest of this paper.
@@ 3.1 Mixed-Integer Linear Programming with Fixed Depots’ Locations
Assume all orbital elements of depots are fixed. Then, $\phi_{k^{\prime}}$, $\Delta V_{k^{\prime}j}({\bf\it\eta}_{k}^{\prime})$, and $\Delta V_{ik^{\prime}}({\bf\it\eta}_{k}^{\prime})$ are all constants. So, the problem becomes an extension of the VRP, which can be formulated as follows:
@@
$\left[\sum_{p\in D_{s}^{k^{\prime}}}\left(u_{p}-m_{k^{\prime}}^{\text{dry},s}\sum_{j\in T}z_{pj}^{p}\right)+m_{k^{\prime}}^{\text{dry},s}+m_{k^{\prime}}^{\text{dry},d}\right]\phi_{k^{\prime}}\leq m_{\text{max}}^{l}\qquad\forall k^{\prime}\in D_{0}$ (31)
@@ 3.2 Nonlinear Programming with Fixed Routing Solution
$m^{p}({\bf\it\eta})$ $=\left(\left(\left(m^{\text{dry},s}\exp\left[\frac{\Delta V_{n+1}}{g_{0}I_{sp}^{s}}\right]+m_{n}^{L}\right)\exp\left[\frac{\Delta V_{n}}{g_{0}I_{sp}^{s}}\right]+m_{n-1}^{L}\right)\exp\left[\frac{\Delta V_{n-1}}{g_{0}I_{sp}^{s}}\right]+\cdots+m_{1}^{L}\right)\exp\left[\frac{\Delta V_{1}}{g_{0}I_{sp}^{s}}\right]-m^{\text{dry},s}-\sum_{i=1}^{n}m^{L}_{i}$ (32)
@@
where $A=m^{\text{dry,s}}\exp\left[\frac{\sum_{i=2}^{n}\Delta V_{i}}{g_{0}I_{sp}^{s}}\right]$, $B=\sum_{i=1}^{n}m_{i}^{L}\exp\left[\frac{\sum_{j=2}^{i}\Delta V_{j}}{g_{0}I_{sp}^{s}}\right]$, and $C=m^{\text{dry,s}}+\sum_{i=1}^{n}m_{i}^{L}$, which are all constants.
@@
$\min_{{\bf\it\eta}}J({\bf\it\eta})=\min_{{\bf\it\eta}_{k^{\prime}}}\sum_{k^{\prime}\in D_{0}}\sum_{r\in R_{k^{\prime}}}m_{r,k}^{p}({\bf\it\eta}_{k^{\prime}})\phi_{k^{\prime}}({\bf\it\eta}_{k^{\prime}})$ (35)
@@
$r_{\text{min}}$ $\leq a_{k^{\prime}}\qquad\forall k^{\prime}\in D_{0}$ (36)
@@
The partial derivative of $J$ with respect to ${\bf\it\eta}_{k^{\prime}}$ is given as:
$\frac{\partial J}{\partial{\bf\it\eta}_{k^{\prime}}}=\sum_{r\in R_{k^{\prime}}}\left(\frac{\partial m_{r,k}^{p}}{\partial{\bf\it\eta}_{k^{\prime}}}\phi_{k^{\prime}}+m_{r,k}^{p}\frac{\partial\phi_{k^{\prime}}}{\partial{\bf\it\eta}_{k^{\prime}}}\right)\qquad\forall k^{\prime}\in D_{0}$ (37)
For notational convenience, the subscript $k^{\prime}$ is omitted without loss of generality because the derivative that will be derived below can be applied to any $k^{\prime}$, and they are independent of each other. Mathematically,
@@
Also, we use the canonical unit, i.e., $\mu=1~\text{DU}^{3}/\text{TU}^{2}$ and $1~\text{DU}$ is the orbital radius of the constellation (e.g., 26,560 km for GPS constellation). Then, the derivative of $\phi(a)$ with respect to $a$ is as follows:
@@
Now, consider the derivative of $m^{p}_{r,k}$ with respect to $a_{k}$. Similarly, for notational convenience, the subscripts $r$ and $k$ are omitted without loss of generality because the derivative that will be derived below can be applied to any route $r$ of depot $k$. From Eq. (6), $\Delta V_{1}$ and $\Delta V_{n+1}$ are the function of depot’s orbital elements, $\bf\it\eta$}, given as follows.
@@
[…]
@@
[…]
@@
$\partialderivative{\Delta\theta_{1}}{i}$ $=\begin{cases}-\frac{1}{\sqrt{1-\cos^{2}\Delta\theta_{1}}}\quantity(\cos i\sin i_{1}\cos\quantity(\Omega-\Omega_{1})-\sin i\cos i_{1}),&\Delta\theta_{1}\leq 2\\ 0,&\text{otherwise}\end{cases}$ (50)
$\partialderivative{\Delta\theta_{1}}{\Omega}$ $=\begin{cases}\frac{1}{\sqrt{1-\cos^{2}\Delta\theta_{1}}}\sin i\sin i_{1}\sin\quantity(\Omega-\Omega_{1}),&\Delta\theta_{1}\leq 2\\ 0,&\text{otherwise}\end{cases}$ (51)
$\partialderivative{\Delta\theta_{n+1}}{i}$ $=\begin{cases}-\frac{1}{\sqrt{1-\cos^{2}\Delta\theta_{n+1}}}\quantity(\cos i\sin i_{n+1}\cos\quantity(\Omega-\Omega_{n+1})-\sin i\cos i_{n+1}),&\Delta\theta_{n+1}\leq 2\\ 0,&\text{otherwise}\end{cases}$ (52)
$\partialderivative{\Delta\theta_{n+1}}{\Omega}$ $=\begin{cases}\frac{1}{\sqrt{1-\cos^{2}\Delta\theta_{n+1}}}\sin i\sin i_{n+1}\sin\quantity(\Omega-\Omega_{n+1}),&\Delta\theta_{n+1}\leq 2\\ 0,&\text{otherwise}\end{cases}$ (53)
Note that $i_{1}$$i_{n+1}$ and $\Omega_{1}$$\Omega_{n+1}$ represent the inclination and RAAN of the first and last satellite orbits, respectively. Therefore, the first derivatives of the objective function can be calculated analytically, ensuring fast computation and accuracy for the L-BFGS-B algorithm.
The computed gradients are directly passed to the L-BFGS-B optimizer, which uses them to approximate the inverse Hessian and update the decision variables (i.e., the orbital elements ${\bf\it\eta}_{k^{\prime}}$ of each depot) within the given bounds. In each iteration, the optimizer evaluates both the objective function $J({\bf\it\eta})$ and its gradient $\partial J/\partial{\bf\it\eta}$ to decide on the search direction and step size, based on a limited-memory quasi-Newton method. Since the gradients are available analytically in our problem, we avoid numerical differentiation, which helps reduce the number of function evaluations and prevents issues like truncation error or subtractive cancellation.